# Visuomotor Policy Learning via Coarse-to-Fine Autoregressive Prediction

### v2

@@ 6 Conclusion
In this work, we introduce Coarse-to-Fine Autoregressive Policy (CARP), a novel paradigm for robotic visuomotor policy learning that combines the efficiency of autoregressive modeling (AM) with the high performance of diffusion modeling (DM). CARP incorporates: 1) multi-scale action tokenization to capture global structure and temporal locality, addressing AM’s limitations in long-term dependency; 2) coarse-to-fine autoregressive prediction that refines actions from high-level intentions to detailed execution, achieving DM-like performance with AM-level efficiency through latent space prediction and relaxed Markovian constraints. The comprehensive evaluations, from simulation to real-world, demonstrate CARP’s effectiveness in balancing high performance, efficiency, and flexibility.
We hope this work will inspire further exploration into next-generation policy learning by leveraging GPT-style autoregressive models, advancing a more unified perspective on current generative modeling techniques (see Appx.A).
Supplementary Material
@@ Appendix A Limitations and Future Work
In this work, we propose CARP, a next-generation paradigm for robotic visuomotor policy learning, which effectively balances the long-standing trade-off between high performance and high efficiency seen in previous autoregressive modeling (AM) and diffusion modeling (DM) approaches. Despite these advancements, there remain several limitations and opportunities for improvement in future research.
@@
First, the model structure of CARP can be further optimized for simplicity. Currently, CARP employs a two-stage design, where the first stage utilizes separate multi-scale action VQVAE modules for each action dimension to address their orthogonality. Future work could focus on developing a unified one-stage method that integrates multi-scale tokenization with the coarse-to-fine prediction process, resulting in a more efficient and streamlined framework without compromising performance.
Second, the multimodal capabilities of CARP remain underdeveloped. To mitigate the unimodality issue inherent in traditional autoregressive policies trained with MSE loss, CARP adopts a cross-entropy loss, which retains the potential for multimodality. Compared to the Diffusion Policy’s multimodal ability [11] likely coming from DDPM’s integration over a Stochastic Differential Equation [19], CARP’s approach is relatively straightforward. Further enhancements are needed to fully realize and leverage its multimodal potential.
@@
Third, CARP’s adoption of the GPT-style paradigm opens up promising yet unexplored possibilities. Beyond the flexibility already demonstrated, the contextual understanding capabilities inherent in GPT-style architectures [44] suggest that CARP could be extended to support multimodal inputs [3, 45] like tactile and auditory information and address robotic tasks requiring long-term dependency reasoning [1]. Additionally, its potential for generalization in few-shot and zero-shot learning scenarios [7] represents a particularly valuable direction for future exploration.
@@
Finally, but not exhaustively, the scaling properties of CARP represent a promising frontier. The scaling laws established in existing GPT-style models [25] could be seamlessly applied to CARP, suggesting that increasing network capacity and leveraging larger pre-training datasets could lead to substantial performance gains. Furthermore, recent advances in Vision-Language-Action (VLA) [6, 27, 31] models present an opportunity to integrate CARP into such frameworks, which could further validate its scalability and capabilities.
@@
Figure 10: Coarse-to-Fine Autoregressive Inference. During inference, we leverage kv-caching to enable coarse-to-fine prediction without the need for causal masks. The full set of token maps, $\boldsymbol{r}_{1:K}$, is collectively decoded by the action VQVAE into executable actions for the robotic arm.
Appendix B Definition of Autoregressive Policy
Figure 11: Conventional Autoregressive Policy. In reinforcement learning, conventional autoregressive policies generate action tokens sequentially, where each token is predicted based on the previously generated tokens. This differs from the action chunking prediction shown in Fig. 2(a).
@@ Appendix B Definition of Autoregressive Policy
Autoregressive policies naturally capitalize on the efficiency and flexibility of autoregressive models. Initial works from a reinforcement learning perspective applied models like Transformers to predict the next action using states or rewards as inputs [10, 21], as shown in Fig. 11 and formalized as:
@@
More recently, the concept of action chunking [28], derived from neuroscience, has demonstrated notable benefits for imitation learning [52, 12, 70, 30]. In action chunking, individual actions are grouped and executed as cohesive units, leading to improved efficiency in storage and execution, as depicted in Fig. 2(a). This paradigm extends the capabilities of GPT-style decoders by modifying them to generate chunks of actions in one forward pass, replacing the traditional single-step autoregressive operation with a multi-token, pseudo-autoregressive process. The action generation process for this paradigm is described as:
@@
Given the significant performance improvements enabled by action chunking, we adopt this multi-token, one-forward-pass framework throughout the article and experiments when referring to autoregressive policies (AP).
Appendix C Implementation Consistent with Diffusion Policy
We adopt similar experimental settings with 1 or 2 observations, a prediction horizon of 16, and an executable action length of 8, following the standard setup used in Diffusion Policy [11].
@@ Appendix C Implementation Consistent with Diffusion Policy
It is important to note that our classical formulation introduces a minor discrepancy in the horizon definition compared to the implementation of Diffusion Policy [11]. Specifically, in our formulation, the horizon $H$ encompasses the past observed steps, meaning that the index of the current next predicted action is $O$, rather than 0. In contrast, as outlined in formulation Eq. 1, the horizon $H$ does not include past observations, with the first prediction step corresponding to the next time step. While the rationale behind this design remains unclear due to limitations in the author’s understanding, we retain the horizon definition introduced by Diffusion Policy [11] to ensure consistency in our experimental comparisons.

### v3

@@ 6 Conclusion
In this work, we introduce Coarse-to-Fine Autoregressive Policy (CARP), a novel paradigm for robotic visuomotor policy learning that combines the efficiency of autoregressive modeling (AM) with the high performance of diffusion modeling (DM). CARP incorporates: 1) multi-scale action tokenization to capture global structure and temporal locality, addressing AM’s limitations in long-term dependency; 2) coarse-to-fine autoregressive prediction that refines actions from high-level intentions to detailed execution, achieving DM-like performance with AM-level efficiency through latent space prediction and relaxed Markovian constraints. The comprehensive evaluations from single- to multi-task, simulation to real-world, demonstrate CARP’s effectiveness in balancing high performance, efficiency, and flexibility.
We hope this work inspires further exploration of next-generation GPT-style autoregressive models for policy learning, fostering a unified perspective on current generative modeling techniques (see Suppl. A for further discussion).
Acknowledgement
This work was supported by the National Science and Technology Innovation 2030 - Major Project (Grant No. 2022ZD0208800)
@@ Appendix A Limitations and Future Work
In this work, we propose CARP, a next-generation paradigm for robotic visuomotor policy learning, which effectively balances the long-standing trade-off between high performance and high inference efficiency seen in previous autoregressive modeling (AM) and diffusion modeling (DM) approaches. Despite these advancements, there remain several limitations and opportunities for improvement in near-future research.
@@
First, the architectural design of CARP can be further optimized for simplicity. Currently, CARP employs a two-stage design, where the first stage utilizes separate multi-scale action VQVAE modules for each action dimension to address their orthogonality. A promising direction for future work could focus on developing a unified one-stage method that integrates multi-scale tokenization with the coarse-to-fine prediction process, resulting in a more efficient and streamlined framework without compromising performance.
Second, CARP’s multimodal capacity has not yet been fully explored or leveraged. To address the inherent unimodality of conventional autoregressive policies trained with MSE loss, CARP employs a Cross-Entropy objective that preserves the potential for multi-modal predictions. Compared to the Diffusion Policy [12]’s ability to model multi-modality via DDPM’s integration over stochastic differential equations [21], CARP adopts a more direct yet effective alternative that achieves comparable multimodal expressiveness. Nevertheless, the role of multi-modality in visuomotor policy learning remains underexplored. Many current benchmark tasks either do not require diverse output distributions or tend to induce overfitting to a single prediction path. Future research should investigate the necessity of multi-modal reasoning in robotic decision-making and further harness CARP’s capacity to model action diversity.
@@
Third, CARP’s adoption of the GPT-style paradigm opens up promising yet unexplored possibilities. Beyond the flexibility already demonstrated, the contextual understanding capabilities inherent in GPT-style architectures [46] suggest that CARP could be extended to support multi-modal inputs [3, 47] like tactile and auditory information and address robotic tasks requiring long-term dependency reasoning [1]. Moreover, its inherent capacity for in-context learning suggests strong potential for generalization under few-shot and zero-shot learning settings [8], making it a compelling foundation for more adaptive and versatile visuomotor policies.
@@
Finally, but not exhaustively, the scaling potential of CARP presents a promising avenue for future exploration. The scaling laws established in existing GPT-style models [27] could be seamlessly applied to CARP, suggesting that increasing model capacity and leveraging larger pre-training datasets could lead to substantial performance gains. Furthermore, recent advances in Vision-Language-Action (VLA) [7, 29, 33] models present a promising opportunity to integrate CARP into such frameworks. Such integration could further demonstrate CARP’s scalability and its potential for general-purpose embodied intelligence.
Appendix B Coarse-to-Fine Inference
Unlike the training process, the inference process predicts token maps of the action sequence across different scales in an autoregressive next-scale, coarse-to-fine manner without teacher forcing, as illustrated in Fig. 8. Additionally, kv-caching is employed to eliminate redundant computations.
@@ Appendix B Coarse-to-Fine Inference
Figure 8: Coarse-to-Fine Autoregressive Inference. During inference, we leverage kv-caching to enable coarse-to-fine prediction without the need for causal masks. The full set of token maps, $\boldsymbol{r}_{1:K}$, is collectively decoded by the action multi-scale VQVAE into executable actions for the robotic arm.
Appendix C Definition of Autoregressive Policy
@@
Autoregressive policies naturally capitalize on the efficiency and flexibility of autoregressive models. Initial works from a reinforcement learning perspective applied models like Transformers to predict the next action using states or rewards as inputs [11, 23], as shown in Fig. 9 and formalized as
@@
Figure 9: Conventional Autoregressive Policy. In reinforcement learning, conventional autoregressive policies generate action tokens sequentially, where each token is predicted based on the previously generated tokens. This differs from the action chunking prediction (see Sec. 2.2 of the main paper).
More recently, the concept of action chunking [30], derived from neuroscience, has demonstrated notable benefits for imitation learning  [54, 13, 72, 32]. In action chunking, individual actions are grouped and executed as cohesive units, leading to improved efficiency in storage and execution, as depicted in Fig. 2a of the main paper. This paradigm extends the capabilities of GPT-style decoders by modifying them to generate chunks of actions in one forward pass, replacing the traditional single-step autoregressive operation with a multi-token, pseudo-autoregressive process. The action generation process for this paradigm is described as
@@
Given the significant performance improvements enabled by action chunking, we adopt this multi-token, one-forward-pass framework throughout the article and experiments when referring to Autoregressive Policy (AP).
Appendix D Efficiency Concerns
Efficiency, as discussed throughout this paper, specifically refers to inference efficiency—the ability of CARP to generate actions significantly faster than diffusion-based policies during deployment. While efficiency can be examined from various perspectives, we focus on three key aspects relevant to CARP: inference efficiency, training efficiency, and data efficiency, each of which is analyzed in detail below.
Task CARP DP
Predict Decode
Can 2.279 s 0.639 s 34.79 s
Square 2.597 s 0.679 s 35.62 s
Table 5: Inference Efficiency Comparison. We report the time consumption for the coarse-to-fine prediction phase and the subsequent action decoding phase over 400 timesteps of action generation. The results indicate that in CARP, the majority of inference time is allocated to the prediction step, whereas the decoding process is completed within a short duration.
Inference Efficiency. We analyze the inference efficiency of CARP’s two-stage process. During inference, CARP first predicts action tokens in a coarse-to-fine manner, followed by a single forward pass to decode all token maps into executable actions. Since token maps are collected during the prediction phase, and decoding requires only a single forward computation, the majority of the computational cost is incurred during prediction, while action decoding remains relatively lightweight. This is empirically validated by the results in Tab. 5. Compared to DP, CARP achieves significantly faster inference by eliminating the iterative denoising steps required by diffusion-based policies, instead directly predicting actions as a low-dimensional generation problem.
Training Efficiency. We analyze training efficiency through convergence behavior and time consumption.
While training convergence depends on task complexity and hyperparameter configurations, both CARP and DP exhibit stable learning dynamics under our respective settings. To accommodate the architectural differences of CARP, we employ slightly different training configurations from those used for DP. As shown in Fig. 10, under our experimental settings, both DP and CARP achieve good convergence within the same number of training epochs. While convergence speed may differ due to structural and training differences, it does not inherently indicate superiority in model design. Instead, both CARP and DP demonstrate reliable training behavior under their respective settings.
Figure 10: Training Efficiency on Convergence Analysis. With different training configurations, both DP and CARP converge effectively within 4000 epochs in the state-based Square task.
Figure 11: Training Efficiency on Time Consumption. Comparison of training time on the Square task. For CARP, we separately report the tokenizer training time (with 10 VQ-VAEs, one per action dimension) and policy learning time.
In terms of wall-clock training time, a comparison is provided in Fig. 11. The total training cost of CARP is comparable to that of DP when considering the policy learning stage (CARP-TF) alone. Although CARP introduces an additional tokenizer pretraining phase (CARP-VQ)—where separate VQ-VAEs are trained for each action dimension—the cost is amortized across tasks and environments. In particular, when trained on sufficiently diverse data (e.g., multi-task settings in both simulation and the real world), the tokenizers become reusable, substantially reducing the overall training burden in practical deployments.
Data Efficiency. We further assess the data efficiency of CARP by evaluating its performance under varying amounts of training data. Specifically, we investigate whether CARP can maintain strong performance when trained with limited trajectories, indicating robustness to data scarcity. As shown in Fig. 12, we compare CARP with baseline policies (following implementations from [12]) on the state-based Square task, using training datasets ranging from 200 to 30 trajectories. CARP consistently outperforms the baselines across all data regimes, demonstrating its superior data efficiency and reduced reliance on large-scale datasets.
Figure 12: Data Efficiency Analysis. Performance comparison under varying training dataset sizes on the Square task. Each policy is trained following its official best-practice settings. CARP consistently outperforms the baselines at all data scales.
Appendix E Comparative Analysis of CARP and AR.
CARP outperforms traditional autoregressive (AR) models in terms of success rate while maintaining high computational efficiency. CARP adopts a straightforward action tokenizer and leverages a GPT-style transformer for prediction, similar to standard AR policies. However, instead of conventional next-token prediction, CARP introduces a paradigm shift towards next-scale prediction. Despite these seemingly minor modifications, CARP achieves substantial performance gains. In this section, we analyze the key factors contributing to this improvement.
Temporal Locality. CARP encodes action sequences into a latent space in its first stage. Specifically, it employs 1D convolution to explicitly capture the local correlations within actions, facilitating a more effective learning of temporal dependencies—something that step-by-step action modeling struggles to achieve. As depicted in the magnified region of Fig. 14, encoding actions into a latent space enhances the smoothness of predictions while simultaneously denoising raw actions. This encoding process enables the model to capture similarities and overarching trends across contiguous actions, leveraging temporal locality to its advantage. While recent work on action chunking [72] has highlighted the significance of temporal locality, existing next-token prediction models still suffer from weakened action dependencies due to their traditional independent action output mechanisms.
Global Structure. CARP represents action sequences across multiple scales and predicts actions in a coarse-to-fine manner. The coarser scales compress the sequence using fewer tokens, promoting the learning of global action patterns. This hierarchical representation explicitly models the overall structure of action sequences—an aspect that traditional unidirectional next-token prediction struggles to capture. By progressively refining actions from high-level to low-level representations, CARP enhances sequence stability and mitigates the risk of producing erratic, inconsistent motions, leading to more precise execution.
Action Scalability. Similar to the approach used in Diffusion Policy, encoding action sequences improves the scalability of action generation. By encoding action sequences, CARP enables flexible adjustments to sequence length with minimal modifications to the model architecture, offering greater adaptability across different tasks and environments.
Figure 13: Raw vs. Reconstructed Actions (Jointly Trained VQ-VAE). Flattening and jointly encoding all action dimensions into a single VQ-VAE leads to poor reconstruction, as shown by the large discrepancy between raw (blue) and reconstructed (red) trajectories. This highlights the limitation of naive joint encoding.
Figure 14: Comparison of Raw and Reconstructed Actions. Comparison across 10 action dimensions in the Coffee and Stack Three tasks. Reconstructed actions (red) closely align with raw signals (blue), preserving structural patterns while smoothing the sequences and filtering out noise (see magnified region), highlighting the effectiveness of our multi-scale tokenization.
Precision in Encoding. Adapting the multi-scale VQ-VAE to action space necessitates careful architectural design. Rather than flattening the action trajectory into an image-like structure and training a single joint VQ-VAE—which leads to unstable and less interpretable tokens (Fig. 13)—we instead encode each action dimension independently, using a dedicated VQ-VAE per dimension. As illustrated in Fig. 14, the multi-scale tokenization process ensures that generated action sequences closely match the raw inputs, exhibiting nearly identical trajectory lines while yielding smoother motions. This demonstrates that our action tokenization approach effectively preserves the fidelity of original action sequences. Moreover, the enhanced success rates observed in the experiments presented in the main paper further validate the accuracy and effectiveness of CARP’s design.
Appendix F Additional Baselines Comparison
In addition to the baseline policies discussed in the main paper, we further compare CARP with enhanced versions of each category: VQ-BET [32], an improved variant of BeT [54], and Consistency Policy [44], which reduces sampling steps to improve inference efficiency. Both are implemented using their official codebases with recommended settings. As shown in Tab. 6, CARP achieves consistently higher success rates across all tasks, while maintaining competitive inference time. These results highlight CARP as a promising design that combines strong task performance with high inference efficiency.
Policy p1 p2 p3 p4 Inf.T(s)
ConsisP 0.99 0.96 0.95 0.93 2.31
VQ-BeT 0.96 0.92 0.87 0.71 1.48
CARP 1.00 1.00 0.98 0.98 2.01
Push-T Inf.T(s)
0.80 2.93
0.72 1.70
0.88 2.66
Table 6: State-Based Kitchen and Push-T Results. We compare CARP with VQ-BET [32] and Consistency Policy [44] under identical settings. CARP consistently outperforms both baselines in success rate, while offering competitive inference times.
Appendix G Failure Analysis
In this section, we analyze common failure cases observed during experiments with both the diffusion-based policies and our proposed CARP.
Accident Recovery. A notable failure mode is the inability to recover from disturbances, as illustrated in Fig. 15. When tools are accidentally knocked over due to suboptimal action trajectories, the model struggles to generate appropriate recovery behaviors. This limitation arises because the policy is trained purely by imitating expert demonstrations, which do not account for such out-of-distribution failure scenarios. Addressing this issue requires further incorporation mechanisms for failure detection and recovery.
Hesitant Movements. Another common failure case involves the generation of jerky or oscillatory movements when the robot encounters two similar situations with only slight visual differences across consecutive timesteps, as shown in the first row of Fig. 16. This issue arises because the policy conditions its predictions on only the previous one or two observations, potentially overlooking long-term historical context. When faced with multiple plausible action choices under limited observations, the policy may produce ambiguous actions, leading to hesitation. Consequently, these hesitant movements can prevent the robot from meeting the success criteria, as shown in the bottom-right of Fig. 16.
Figure 15: Accident Recovery in the Square Task. When tools are accidentally knocked over, the robot struggles to recover due to its reliance on imitation learning from expert demonstrations, which lack exposure to such out-of-distribution failure cases.
Figure 16: Hesitant Movements in the Mug Task. During task execution, the robot may encounter visually similar observations at different timesteps. Without leveraging long-term historical context, the policy may misinterpret these similarities and generate ambiguous actions, resulting in hesitant (jerky) movements. In this failure case, after successfully closing the drawer, the policy perceives the scene as similar to the initial step and erroneously attempts to reopen it. This cycle of opening and closing continues indefinitely, leading to task failure.
Appendix H Experiment Implementation Details
Here, we provide implementation details for the main experiments presented in the paper.
Single-Task Simulation Experiment. For baseline models, we follow the same implementation and training configurations provided by Diffusion Policy. For all state-based experiments, including the Kitchen and Push-T tasks, we uniformly set the observation horizon $O=2$ and the prediction horizon $H=16$ across all models. For image-based experiments, we set $O=1$ and $H=16$ for better transferability to real-world scenarios. As per the benchmark, only the first 8 actions in the prediction horizon are executed, starting from the current step (see Suppl. I for further discussion). For CARP, we first train an action VQVAE model (see Sec. 3.1 of the main paper) following [31], using $V=512$, $C=8$, a batch size of 256, and 300 epochs per task. Given a horizon $H=16$, we design multi-scale representations with scales of $1,2,3$ and $4$ to capture coarse-to-fine information across the action sequence. We then train an autoregressive GPT-2 style, decoder-only transformer (see Sec. 3.2 of the main paper) , based on [60], using the same training settings as the benchmark, with a batch size of 256 for state-based experiments (4000 epochs) and a batch size of 64 for image-based experiments (3000 epochs). We use Cross-Entropy loss during training, which preserves the model’s sampling capability. During inference, we typically select the token with the highest probability at each scale. However, to visualize multi-modal behavior in the Push-T task, we sample the top-$k$ tokens at each scale (with k=3k{=}3), allowing for diverse predictions with controlled randomness.
Multi-Task Simulation Experiment. To enable multi-task generalization, CARP augments the single-task formulation by introducing a learnable 3-dimensional task embedding for each task. These embeddings, retrieved based on task indices, are concatenated with the observation sequence $\boldsymbol{s}$ and act as additional conditional inputs to the policy. In our experiments involving 8 tasks, this corresponds to an $8\times 3$ embedding matrix. We also use a moderately deeper decoder-only transformer in GPT-2 style. CARP is trained with a batch size of 512 for 200 epochs on an A100 GPU. Baseline models follow the same training settings as SDP [63]. This minimal modification enables CARP to adapt to multi-task learning seamlessly.
Real-World Experiment. For both baselines and CARP, the input consists of current visual observations from the wrist and scene cameras (resolution: $120\times 160$), as well as proprioceptive data from the robotic arm. We execute 8 predicted actions out of a horizon of 16 predictions. We train the diffusion policy for 3000 epochs with a batch size of 64. For CARP, we use the same visual policy structure as in the simulation tasks, training the multi-scale action tokenizer for 300 epochs with a batch size of 256, and the coarse-to-fine transformer for 3000 epochs with a batch size of 64.
Appendix I Consistent with Diffusion Policy
@@
We adopt similar experimental settings with 1 or 2 observations, a prediction horizon of 16, and an executable action length of 8, following the standard setup used in Diffusion Policy (DP) [12]. It is important to note that our classical formulation introduces a minor discrepancy in the horizon definition compared to the implementation of DP. Specifically, in DP’s experimental setting, the horizon $H$ encompasses the past observed steps, meaning that the index of the current next predicted action is $O$, rather than 0. In contrast, as outlined in the formulation of Eq. (1) in the main paper, the horizon $H$ does not include past observations, with the first prediction step corresponding to the next time step. While the rationale behind this design remains unclear due to limitations in the author’s understanding, we retain the horizon definition introduced by Diffusion Policy (DP) [12] to ensure consistency in our experimental comparisons.
Appendix J Ablation Study on the Number of Scales
Figure 17: Ablation Study on $K$. We evaluate the performance of CARP across three tasks using six different scale configurations. Results indicate that when the number of scales exceeds 4, the model achieves optimal performance. Considering both model efficiency and performance, we set $K=4$ in all experiments throughout the paper.
To maintain consistency with Diffusion Policy, we set the action prediction horizon $H$ to 16 across all tasks. Given $H=16$, we adopt $K=4$ for all experiments. To further investigate the impact of $K$, we conduct an ablation study by varying $K$ from 1 to 6 on three representative tasks: Can, Square, and Kitchen (which requires executing four consecutive subtasks, thus we report success rate based on the final subtask, denoted as p4). All other experimental settings remain unchanged.
For each chosen number of scales $K$, the token map sizes at each scale level are defined as $1,\dots,K$}. Notably, when $K=5$ and $K=6$, the scales slightly exceed the default feature map size. To ensure fair evaluation, we appropriately expand the feature map size to accommodate these settings.
As shown in Fig. 17, using fewer scales leads to insufficient action tokenization, resulting in less precise predictions. When $K=4$, the policy effectively meets task requirements. Further increasing $K$ results in stable performance with negligible fluctuations, indicating that the policy has likely reached its peak performance for the given tasks, with additional scaling providing little to no further benefit.
Appendix K Additional Real-World Experiment
We provide further visualization of the real-world experiments presented in the main paper. As shown in Fig. 18, CARP generates smooth and successful trajectories for the Cup and Bowl tasks, with temporal progression illustrated from left to right.
Figure 18: Visualization of CAPR on Real-World Tasks. CARP generates smooth and successful trajectories on the Cup and Bowl tasks, progressing from left to right.
Beyond the real-world evaluations on a UR5e robot arm, we further deploy CARP on a distinct robotic embodiment: a 7-DoF Franka Emika Panda arm. For this, we adopt the challenging FurnitureBench benchmark [20], which comprises long-horizon, contact-rich manipulation tasks (e.g., pick, place, insert, screw, and flip), with episodes spanning up to 1000 steps (700 steps for One_Leg). A corresponding standard simulator is also provided, as shown in Fig. 19.
We first evaluate CARP and Diffusion Policy (DP) in simulation on three tasks, followed by real-world deployment of the One_Leg task. All experiments are conducted in the state-based setting. To bridge the sim-to-real gap during real-world deployment, 6-DoF object poses are estimated using AprilTags provided by FurnitureBench [20], enabling consistent state-based policy execution.
For simulation, we use 200 trajectories per task from [5]. DP is trained using its official implementation [12] with 100 DDPM denoising steps. CARP follows the same state-based, single-task setting. We evaluate success rates using 1024 rollouts per task for statistical stability. As shown in Tab. 7, CARP achieves competitive performance while offering significantly lower inference time and fewer parameters.
Figure 19: FurnitureBench Tasks for Evaluation. All three tasks are first evaluated in simulation. A corresponding real-world environment is then constructed to assess the performance in real-world.
Policy One_Leg Round_Table Lamp Inf.Time $\downarrow$ Params $\downarrow$
DP-C 39.62% 5.76% 3.91% 74.05s 66.06M
CARP 43.75% 6.25% 4.30% 6.29s 2.54M
Table 7: Simulation Results on FurnitureBench (State-Based). All policies are trained under identical single-task settings. Compared to DP-C [12] (with 100 DDPM denoising steps), CARP achieves comparable success rates while offering significantly lower inference time and parameter count.
We further evaluate CARP on the real-world One_Leg task using 40 expert demonstrations collected via a 3D SpaceMouse (left panel of Fig. 20). The task involves complex rotations and contact-rich interactions, as illustrated in the right panel of Fig. 20. CARP achieves higher success rates across key stages, especially in precision-critical steps such as Insert, demonstrating robust real-world performance.
Appendix L Analysis on Fine-Grained Manipulation
Beyond standard pick-and-place tasks, which are relatively straightforward for robotic manipulation, tasks requiring fine-grained skills have garnered increasing attention. For example, Nut-Assembly and Threading from the MimicGen [41] benchmark demand precise action generation to ensure successful task completion, as illustrated in Fig. 21. In Nut-Assembly, the robot must place a nut onto a designated peg, which is slightly larger in size. Furthermore, the Threading task requires the robot to insert a needle into a small hole on a tripod, a significantly more challenging task due to the minimal margin for error. To evaluate CARP’s fine-grained manipulation capability, we compare it with the state-of-the-art Sparse Diffusion Policy (SDP) [63] in the multi-task setting, and Diffusion Policy (DP) [12] in the single-task setting. We evaluate success rates alongside the mean and variance of the distance between the ideal insertion centers of the fixed structures (peg, tripod) and the centers of the tools (nut, needle) at the moment of first contact. A lower mean distance indicates higher action precision, while a smaller variance reflects the model’s ability to consistently achieve accurate and stable manipulations.
As summarized in Tab. 8 and Tab. 9, our coarse-to-fine autoregressive prediction framework demonstrates strong performance in fine-grained tasks, achieving competitive results comparable to diffusion-based policies. Notably, CARP consistently achieves lower mean error and reduced variance across most tasks, regardless of whether in single-task or multi-task settings. Moreover, CARP achieves these results with an inference speed that is 10 times faster than current diffusion-based policies, highlighting its efficiency and effectiveness in fine-grained robotic manipulation.
Appendix M Task Visualizations
In this section, we provide visualizations of the tasks used in our experiments. For the single-task experiment, the corresponding visualizations are presented in Fig. 22. For the multi-task experiment, visualizations are shown in Fig. 23. For the long-horizon, multi-stage Kitchen experiment, we provide visualizations in Fig. 24, along with the sequential execution process in Fig. 25. Finally, for real-world experiments, visualizations are included in Fig. 26.
Figure 20: Real-World Evaluation on the One_Leg Task. The left panel shows the real-world setup, while the right panel illustrates the execution process and stage-wise results (left to right). Success rates at key stages are reported below. CARP produces smoother motions and outperforms baselines in precision-critical phases such as Grasp and Insert.
Figure 21: Visualization of Fine-Grained Manipulation. We evaluate the precision of generated actions by measuring the distance between the ideal and actual contact centers, represented by the dotted yellow line. Experiments are conducted on Nut-Assembly (top row) and Threading (bottom row). The visualization highlights that CARP achieves a comparable level of precision to diffusion-based policies.
Policy Inference Speed $\uparrow$ Nut Assembly Threading
Success $\uparrow$ Mean $\downarrow$ Variance $\downarrow$ Success $\uparrow$ Mean $\downarrow$ Variance $\downarrow$
SDP 8.20 hz 0.54 7.70 2.36 0.70 5.20 1.25
CARP 118.50 hz 0.66 7.30 1.68 0.70 5.50 1.36
Table 8: Fine-Grained Manipulation Study on Multi-Task Setting. CARP demonstrates a high level of precision comparable to diffusion-based policies, as indicated by the similar mean and variance values. Additionally, CARP achieves a significant speed advantage, running over 10 times faster than diffusion-based approaches. This highlights CARP as a superior balance between performance and efficiency.
Policy Inference Speed $\uparrow$ Nut Assembly Threading
Success $\uparrow$ Mean $\downarrow$ Variance $\downarrow$ Success $\uparrow$ Mean $\downarrow$ Variance $\downarrow$
DP-C 10.13 hz 0.80 5.20 1.86 0.88 4.08 1.07
CARP 119.05 hz 0.82 5.12 1.28 0.88 3.92 0.94
Table 9: Fine-Grained Manipulation Study on Single-Task Setting. To eliminate potential underfitting caused by multi-task training, we evaluate fine-grained tasks under single-task settings. CARP achieves success rates on par with DP-C, while offering lower variance and over 10× faster inference. These results highlight CARP’s ability to maintain high precision and stability with greater inference efficiency.
@@
Figure 25: Visualization of the Consecutive Execution in Kitchen Experiment.